# MindFormer: A Transformer Architecture for Multi-Subject Brain Decoding via fMRI

### v1

@@ 4.1 Experimental Results
Experiment Settings. The proposed model is implemented in PyTorch. The single subject model is trained on one NVIDIA RTX 3090 GPU with a 24GB memory, and the multi subject model is trained on one NVIDIA RTX V100 with a 32GB memorys. Across all experiments, the batch size is set to 4 per GPU and the epoch size is 50. The learning rate was set as $3\times 10^{-4}$, and the moment parameters of the AdamW optimization algorithm [45] were set as $\beta_{1}=0.9,\beta_{2}=0.999$.
Dataset. To better understand the task at hand, we illustrate the data used in our study. For all experiments, we used the widely-adopted Natural Scenes Dataset (NSD)  [17], a public fMRI dataset containing high-resolution 7-Tesla fMRI scans of brain responses from eight healthy adult subjects viewing natural scenes from the MS-COCO dataset  [46]. Following common practices  [20, 21, 22, 23, 25], our research primarily uses data from four subjects (subj01, 02, 05, 07) who completed all scan sessions. Specifically, only a subset of data—982 images—was commonly viewed by all four subjects and used as the test set. The remaining data, comprising 8,859 distinct images viewed by each subject, were used as the training set, resulting in 24,980 training samples without averaging across repetitions, similar to the method used by previous research. We utilize the dataset, preprocessed by  [22], which consists of flattened fMRI voxels within the brain volume space corresponding to the “nsdgeneral” brain region. This region, defined by the authors of  [17], includes the subset of voxels that are most responsive to visual stimuli.
Method ##\## Models Low-level High-level PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ EffNet-B $\downarrow$ SwAV $\downarrow$ Takagi et al. [23] 4 -- -- 83.0% 83.0% 76.0% 77.0% -- -- Brain-Diffuser  [21] 4 .254 .356 94.2% 96.2% 87.2% 91.5% .775 .423 MindEye [22] 4 .309 .323 94.7% 97.8% 93.8% 94.1% .645 .367 MindBridge (Single-) [25] 4 .148 .259 86.9% 95.3% 92.2% 94.3% .713 .413 Ours (Single-) 4 .241 .352 93.5% 97.5% 93.5% 93.6% .659 .356 MindBridge  [25] 1 .151 .263 87.7% 95.5% 92.4% 94.7% .712 .418 Ours (Multi-) 1 .243 .345 93.5% 97.6% 94.4% 94.4% .648 .350
Table 1: Quantitative comparison of MindFormer’s decoding performance against other models. Bold: best, underline: second best. The metrics presented are averaged across the data from 4 subjects. Unlike other methods, which generally require a separate model for each subject, our approach and MindBridge consolidate the process into one model. Among them, our approach achieved superior results in all metrics except for the CLIP score.
Figure 5: Performance comparison on limited datasets. The results shown are for Subject 1, evaluated on different dataset sizes: the entire dataset, 4000 samples, 2000 samples, and 500 samples. Two scenarios are compared: (a) single MindFormer scenario, where the MindFormer is trained exclusively on subject 1’s data, and (b) multi MindFormer scenario, where the MindFormer is trained on data from Subjects 1, 2, 5, and 7.
Method # Dataset Low-level High-level PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ EffNet-B $\downarrow$ SwAV $\downarrow$ (a) 500 .160 .308 80.8% 87.5% 77.5% 78.2% .854 .499 (b) 500 .179 .351 87.7% 93.0% 85.5% 85.5% .796 .461 (a) 2000 .221 .344 90.6% 95.6% 88.6% 89.1% .757 .401 (b) 2000 .241 .352 92.4% 96.8% 91.5% 92.1% .719 .409 (a) 4000 .264 .369 93.5% 96.9% 91.6% 92.3% .701 .374 (b) 4000 .227 .327 93.0% 97.5% 93.6% 93.7% .680 .392 (a) all .270 .356 95.2% 98.0% 93.4% 93.6% .660 .354 (b) all .271 .351 95.3% 98.2% 95.1% 94.7% .641 .339
Table 2: Quantitative comparison of the limited dataset size. The above results are from Subject 1. Bold: best. Two scenarios are compared: (a) single MindFormer scenario, where the MindFormer is trained exclusively on subject 1’s data, and (b) multi MindFormer scenario, where the MindFormer is trained on data from Subjects 1, 2, 5, and 7.
Results. To quantitatively compare with other methods, we utilize eight image quality evaluation metrics as outlined in [21]. For assessing low-level properties, we use PixCorr, SSIM  [47], AlexNet(2), and AlexNet(5)  [48]. For evaluating higher-level properties, the metrics of Inception  [49], CLIP  [31], EffNet-B  [50], and SwAV  [51] are employed.
@@
Also, the quantitative metrics presented in Table 1 further support these findings, indicating significant improvements in high-level indicators such as Inception, CLIP, EffNet-B and SwAV. In brain decoding, low-level metrics evaluate the pixel-wise and structural similarity between original and reconstructed images. High-level metrics assess the semantic similarity and how well the reconstructed images capture the meaning and content of the originals. High-level metrics being high indicates that the model effectively captures and reconstructs the complex, abstract features of the stimulus, leading to more meaningful and contextually accurate representations. These results underscore the effectiveness of our model in accurately decoding and reconstructing visual experiences from neural data, thereby validating our approach and its contributions to the field. Not only does our model achieve high scores on high-level metrics, but it also consistently outperforms other models trained on data from multiple subjects, such as MindBridge, as demonstrated in Table 1. This table illustrates that our model attains superior scores across a range of metrics when compared to MindBridge, which is also designed to handle data from multiple subjects within a single model. These results suggest that our model is more effective in multi-subject brain decoding tasks, thereby implying its potential for broader adoption and application in the field. By excelling in both high-level semantic representation and overall performance, our model demonstrates a significant advancement in the ability to decode and reconstruct brain activity from multiple individuals within a unified framework.
4.2 Ablation Study
Exploiting multiple subject data set.
Given the limited training data, obtaining a sufficiently large fMRI dataset from new subjects is a challenging task. Therefore, it is crucial to achieve high-quality result images even with a small amount of data. To investigate this, we perform ablation study using single MindFormer and multi MindFormer experimental setups. In single MindFormer scenario, MindFormer trains only the dataset of Subject 1. In multi MindFormer scenario, the training process includes not only subject 1 but also subjects 2, 5, and 7. Data from each of these subjects is processed through the MindFormer framework, allowing the model to learn from a varied set of neural patterns. This approach ensures that the model can generalize more effectively across different individuals, thereby enhancing its decoding accuracy and robustness. By incorporating multiple subjects, the MindFormer can leverage the collective information, leading to improved performance in brain decoding tasks.
Figure 5 and Table 2 presents the reconstructed images and metric results for subject 1 across different dataset sizes: the entire dataset, 4000 samples, 2000 samples, and 500 samples. Overall, the results indicate that the multi MindFormer approach consistently outperforms the single MindFormer. Notably, the results obtained using only 2000 samples demonstrate that multi MindFormer, which is our method, can effectively reconstruct high-quality images even with a limited amount of data. Although the performance degraded with 500 samples, multi MinFormer scenario still outperforms the single MindFormer. Specifically, Figure 5 shows that as the dataset size decreases, the single MindFormer struggles to preserve the semantic aspects of the stimulus image, whereas the multi MindFormer maintains this semantic fidelity well. This highlights the robustness and efficiency of multi MindFormer in leveraging small datasets to achieve superior image reconstruction.
@@ 4.2 Ablation Study
Table 3: Quantitative comparison of results on the presence or absence of learnable subject tokens (ST). Bold: best. The inclusion of subject tokens significantly improved the model’s accuracy, as evidenced by the higher scores across various evaluation metrics. This demonstrates that learnable subject tokens play a crucial role in enhancing the precision and reliability of the brain decoding process.
Importance of subject tokens.
@@
Using learnable subject tokens in MindFormer has several positive effects. First, it allows the model to accurately distinguish between inputs from different subjects, ensuring that individual-specific neural patterns are correctly interpreted and processed. This enhances the precision of brain decoding by aligning the model’s understanding with the unique characteristics of each subject’s brain activity. Additionally, the incorporation of learnable subject tokens improves the model’s ability to generalize across multiple subjects, as it can adapt to variability in neural signals while maintaining high performance in decoding tasks. Overall, subject tokens contribute to more reliable and robust decoding outcomes, facilitating better insights into neural representations. Figure 6 and Table 3 provides evidence for these benefits by showing superior performance results when subject tokens are used. Also, by using subject tokens, this unified model approach offers significant advantages in terms of efficiency and scalability, allowing for comprehensive analysis and image reconstruction across different individuals without the need for multiple models.
5 Discussion
@@
The results from our experiments show the efficacy and robustness of the MindFormer model in cross-subject brain decoding. One of the key findings is the significant improvement in image reconstruction accuracy when incorporating subject tokens and utilizing a multi-subject training approach. Notably, even with a limited dataset, the MindFormer demonstrates its capability to reconstruct high-quality images, outperforming models trained on larger datasets. This is particularly important given the challenges associated with obtaining large fMRI datasets from new subjects. The ability to achieve good performance with smaller datasets not only validates the efficiency of the MindFormer but also points to its practical applicability in real-world scenarios where data availability may be constrained. Moreover, the comparison between single-subject and multi-subject training further validates the advantage of leveraging data from multiple subjects. The results consistently indicate that the MindFormer approach, which integrates data from subjects 2, 5, and 7 along with subject 1, yields better performance metrics. This suggests that the model benefits from the additional information provided by the diverse set of neural patterns, enhancing its generalization capabilities and robustness.
@@
Although the MindBridge model [25] is designed to exploit multi-subject fMRI data for training, it relies on an aggregation function to reduce dimensionality, leading to information loss and lower performance. On the other hand, the learnable subject tokens play a pivotal role in our MindFormer framework. By allowing the model to differentiate between neural signals from different subjects, the subject tokens ensure that the individual-specific nuances in brain activity are preserved and accurately decoded. Table 3 demonstrates the performance improvements attributed to the use of subject tokens, highlighting their importance in achieving high-fidelity decoding outcomes. This approach also streamlines the process by enabling the use of a single unified model for multiple subjects, offering significant advantages in terms of efficiency and scalability.
@@ 6 Conclusion
In this paper, we proposed "MindFormer", a new transformer architecture for powerful multi- subject brain decoding framework when combined with Stable Diffusion model. The model successfully surpass the existing multi-subject brain decoding framework. The inclusion of learnable subject tokens contributed to the model’s robustness, enabling it to generalize effectively across different subjects. The proposed method, MindFormer, successfully reconstructs images from fMRI signals with higher accuracy than previous models, thanks to its innovative use of the IP-Adapter and learnable subject tokens. Our contributions include the development of a unified model for multi-subject brain decoding, the creation of a lightweight deep learning model, and the demonstration of strong performance with limited data. Overall, MindFormer provides a new framework for understanding multi-subject brain decoding and common neural patterns. The model’s ability to leverage shared information across subjects while maintaining individual-specific accuracy marks a significant advancement in the field of brain decoding.
Limitation
However, there are several limitations to our approach that must be addressed in future research. First, the current implementation of MindFormer primarily focuses on visual stimuli. Extending this approach to decode more complex cognitive and sensory experiences will require substantial advancements in both model architecture and training methodologies. Another limitation is the computational complexity associated with training much more subjects. Although our approach reduces the parameter count compared to existing models, training these models with over about 10 subjects still require significant computational resources. Future work should aim to optimize the model further to make it more accessible and feasible for broader applications. That said, the small model size could impose limitations on the ability to decode brain activity at the individual subject level. Due to the reduced capacity of the model, it may struggle to capture the unique and complex neural patterns specific to each subject, thereby affecting the accuracy and effectiveness of the decoding process.
Societal Impact
Ethical considerations, such as privacy and consent, are critical in this domain. The ethical implications of visual brain decoding, including privacy concerns and the potential for misuse, must be carefully considered. Collaborative efforts between neuroscientists, computer scientists, and ethicists will be essential in advancing this field responsibly and effectively. As visual brain decoding technologies continue to mature, they hold promise for profound impacts on neuroscience, clinical practice, and our understanding of the human.
@@ A.1 Comparison of Model Parameter Number
Table 4: MindBridge and MindFormer have fewer parameters compared to other models. In particular, the single-MindFormer for one subject has the fewest number of parameters among them. Using models with fewer parameters is a way to reduce computational costs.
@@ A.2 Additional Results of multi MindFormer Model
Figure 7: Additional reconstructed image from human brain activity. All reconstructions are outputs from the our one integrated model, MindFormer.
Figure 8: Additional reconstructed image from human brain activity. All reconstructions are outputs from the our one integrated model, MindFormer.

### v2

@@ 4.1 Experimental Results
To quantitatively compare with other methods, we utilize eight image quality evaluation metrics as outlined in Ozcelik & VanRullen (2023). For assessing low-level properties, we use PixCorr, SSIM (Wang et al. (2004)), AlexNet(2), and AlexNet(5) (Krizhevsky et al. (2012)). For evaluating higher-level properties, the metrics of Inception (Szegedy et al. (2016)), CLIP (Radford et al. (2021)), EffNet-B (Tan & Le (2019)), and SwAV (Caron et al. (2020)) are employed. We compared our model with  Takagi & Nishimoto (2023a), Brain-Diffuser (Ozcelik & VanRullen (2023)), MindEye (Scotti et al. (2024)), and MindBridge (Wang et al. (2024)).
@@
Also, the quantitative metrics presented in Table 1 further support these findings, indicating significant improvements in high-level indicators such as Inception, CLIP, EffNet-B and SwAV. In brain decoding, low-level metrics evaluate the pixel-wise and structural similarity between original and reconstructed images. On the other hand, the high-level metrics assess the semantic similarity and how well the reconstructed images capture the meaning and content of the originals. High-level metrics being high indicates that a model effectively captures and reconstructs the complex, abstract features of the stimulus, leading to more meaningful and contextually accurate representations. Therefore, for the purpose of semantically aligned brain decoding from multi-subject data, high-level metric is more important. Thus, the results in Table 1 confirm the effectiveness of our model in semantically aligned decoding from neural data.
Additionally, not only does our model achieve high scores on high-level metrics, but it also consistently outperforms other models trained on data from multiple subjects, such as MindBridge, in terms of low-level metrics as demonstrated in Table 1. This table illustrates that our model attains superior scores across a range of metrics when compared to MindBridge, which is also designed to handle data from multiple subjects within a single model. These results suggest that our model is more effective in multi-subject fMRI signal alignment and decoding, thereby implying its potential for broader adoption and application in the field. By excelling in both high-level semantic representation and overall performance, our model demonstrates a significant advancement in the ability to decode and reconstruct brain activity from multiple individuals within a unified framework.
@@
Table 2: Quantitative comparison of results on the presence or absence of learnable subject tokens (ST). The inclusion of subject tokens significantly improved the model’s accuracy, as evidenced by the higher scores across various evaluation metrics. This demonstrates that learnable subject tokens play a crucial role in semantic alignment of multi-subject fMRI signal and the reliability of the brain decoding process. Bold: best.
4.2 Ablation Study
@@
Importance of subject tokens. Using learnable subject tokens in MindFormer has several positive effects. First, it allows the model to accurately distinguish between inputs from different subjects, ensuring that individual-specific neural patterns are correctly interpreted and processed. This enhances the precision of brain decoding by aligning the model’s understanding with the unique characteristics of each subject’s brain activity. Additionally, the incorporation of learnable subject tokens improves the model’s ability to generalize across multiple subjects, as it can adapt to variability in neural signals while maintaining high performance in decoding tasks. Overall, subject tokens contribute to more reliable and robust decoding outcomes, facilitating better insights into neural representations. Figure 5 and Table 2 provides evidence for these benefits by showing superior performance results when subject tokens are used. Also, by using subject tokens, this unified model approach offers significant advantages in terms of efficiency and scalability, allowing for comprehensive analysis and image reconstruction across different individuals without the need for multiple models.
Exploiting multiple subject data set. Given the limited training data, obtaining a sufficiently large fMRI dataset from new subjects is a challenging task. Therefore, it is crucial to achieve high-quality result images even with a small amount of data. To investigate this, we perform ablation study using single-subject and multi-subject experimental setups. In single subject scenario, MindFormer trains only the dataset of Subject 1. In multi subject scenario, the training process includes not only subject 1 but also subjects 2, 5, and 7. Data from each of these subjects is processed through the MindFormer framework, allowing the model to learn from a varied set of neural patterns. This ablation study investigates how the model can generalize more effectively across different individuals, thereby enhancing its decoding accuracy and robustness.
Figure 6 and Table 3 presents the reconstructed images and metric results for subject 1 across different dataset sizes: the entire dataset, 4000 samples, 2000 samples, and 500 samples. Overall, the results indicate that the multi-subject approach consistently outperforms the single subject approach. Notably, the results obtained using only 2000 samples demonstrate that multi-subject training, which is our method, can effectively reconstruct high-quality images even with a limited amount of data. For example, even with as low as 500 samples, multi-subject approach still outperforms the single subject approach. Specifically, Figure 6 shows that as the dataset size decreases, the single subject approach struggles to preserve the semantic aspects of the stimulus image, whereas the multi subject approach maintains this semantic fidelity well. This highlights the robustness and efficiency of multi MindFormer in leveraging small datasets to achieve superior image reconstruction. By incorporating multiple subjects, MindFormer can leverage the collective information, leading to improved performance in brain decoding tasks.
Figure 6: Performance comparison on limited datasets. With limited training data from a single subject, our proposed MindFormer can reconstruct natural images more accurately by leveraging knowledge from other subjects. The results shown are for Subject 1, trained on different dataset size. Two scenarios are compared: (a) single subject scenario, where the MindFormer is trained exclusively on subject 1’s data, and (b) multi subject scenario, where the MindFormer is trained on data from subjects 1, 2, 5, and 7.
Method # Dataset Low-level High-level PixCorr $\uparrow$ SSIM $\uparrow$ Alex(2) $\uparrow$ Alex(5) $\uparrow$ Incep $\uparrow$ CLIP $\uparrow$ EffNet-B $\downarrow$ SwAV $\downarrow$ (a) all .270 .356 95.2% 98.0% 93.4% 93.6% .660 .354 (b) all .271 .351 95.3% 98.2% 95.1% 94.7% .641 .339 (a) 4000 .264 .369 93.5% 96.9% 91.6% 92.3% .701 .374 (b) 4000 .227 .327 93.0% 97.5% 93.6% 93.7% .680 .392 (a) 2000 .221 .344 90.6% 95.6% 88.6% 89.1% .757 .401 (b) 2000 .241 .352 92.4% 96.8% 91.5% 92.1% .719 .409 (a) 500 .160 .308 80.8% 87.5% 77.5% 78.2% .854 .499 (b) 500 .179 .351 87.7% 93.0% 85.5% 85.5% .796 .461
Table 3: Quantitative comparison of the limited dataset size. The above results are from Subject 1. Two scenarios are compared: (a) single subject scenario, where the MindFormer is trained exclusively on subject 1’s data, and (b) multi subject scenario, where the MindFormer is trained on data from Subjects 1, 2, 5, and 7. Bold: best.
Figure 7: fMRI-to-Text Model using a pretrained Mindformer and an LLM. For the input of LLM, the output of the MindFormer is mapped to textual embedding using a two-layer MLP.
4.3 fMRI-to-Text Experiments
In order to confirm that the significant improvement of our model is originated from semantically align feature space in MindFormer rather than Stable Diffusion, we additionally perform fMRI-to-text generation experiments by inputting the MindFormer feature as the input of Large Language Model (LLM). This experimental setup is unique as we do not rely on the Stable Diffusion image generator.
Implementation Details. Figure 7 shows the framework of the fMRI-to-Text generation. We employ a simple two-layer MLP to align the image feature embeddings from the pretrained Mindformer to the word embedding space of a LLM. We choose OPT-1.3B model as our LLM. With the Mindformer and OPT-1.3B remain frozen, the two-layer MLP is trained with the language modeling loss between the generated captions and the ground truth COCO captions of subjects 1,2,5 and 7. The entire fMRI-to-Image caption model is trained on NVIDIA A100 with 40GB of memory for 5 epochs with a learning rate of 1e-5 and a batch size of 1.
Results. To quantitatively compare with other methods, we utilize six text quality evaluation metrics. For assessing low-level properties, we use Meteor (Banerjee & Lavie (2005)), Rouge (Lin (2004)), and CIDEr (Vedantam et al. (2015)). For evaluating higher-level properties, the metrics of SPICE (Anderson et al. (2016)), CLIP (Radford et al. (2021)), and Sentence (Reimers (2019)) are employed. We compared our model with SDReconT (Takagi & Nishimoto (2023b)), UniBrain (Mai & Zhang (2023)), BrainCap (Ferrante et al. (2023)), and MindSemantix (Ren et al. (2024)). We have referenced the result values from the MindSemantix. Table 4 demonstrates superior performance in the metrics of Meteor, CIDEr, SPICE, and Sentence, compared to other models. Also, Figure 8 shows the COCO captions (ground-truth) and generated texts from fMRI signals, and our model successfully generated the caption, corresponding to the stimulus image. The comprehensive results indicate that the output embedding of our MindFormer contains the relevant information needed to generate texts effectively.
4.4 Discussion
@@
The results from our experiments show the efficacy and robustness of the MindFormer model in cross-subject brain decoding. One of the key findings is the significant improvement in image reconstruction accuracy when incorporating subject tokens and utilizing a multi-subject training approach. Notably, even with a limited dataset, the MindFormer demonstrates its capability to reconstruct high-quality images, outperforming models trained on larger datasets. This is particularly important given the challenges associated with obtaining large fMRI datasets from new subjects. The ability to achieve good performance with smaller datasets not only validates the efficiency of the MindFormer but also points to its practical applicability in real-world scenarios where data availability may be constrained. Moreover, the comparison between single-subject and multi-subject training further validates the advantage of leveraging data from multiple subjects. The results consistently indicate that the MindFormer approach, which integrates data from subjects 2, 5, and 7 along with subject 1, yields better performance metrics. This suggests that the model benefits from the additional information provided by the diverse set of neural patterns, enhancing its generalization capabilities and robustness. Also, with the output embedding of our MindFormer, LLM can generate the caption corresponding to the stimulus image. It implicitly demonstrates that our Mindformer can be the framework for the fMRI-to-Language generation model.
@@
Although the MindBridge model (Wang et al., 2024) is designed to exploit multi-subject fMRI data for training, it relies on an aggregation function to reduce dimensionality, leading to information loss and lower performance. On the other hand, the learnable subject tokens play a pivotal role in our MindFormer framework. By allowing the model to differentiate between neural signals from different subjects, the subject tokens ensure that the individual-specific nuances in brain activity are preserved and accurately decoded. Table 2 demonstrates the performance improvements attributed to the use of subject tokens, highlighting their importance in achieving high-fidelity decoding outcomes. This approach also streamlines the process by enabling the use of a single unified model for multiple subjects, offering significant advantages in terms of efficiency and scalability.
Figure 8: Results of fMRI-to-text model using MindFormer embedding.
Method Low-level High-level Meteor $\uparrow$ Rouge $\uparrow$ CIDEr $\uparrow$ SPICE $\uparrow$ CLIP $\uparrow$ Sentence $\uparrow$ SDReconT 0.100 0.251 0.138 0.050 0.624 0.280 UniBrain 0.169 0.222 -- -- -- -- BrainCap 0.167 0.407 0.413 0.091 0.705 0.447 MindSemantix 0.190 0.415 0.476 0.125 0.755 0.454 Ours 0.291 0.358 0.634 0.177 0.744 0.485
Table 4: Quantitative results of fMRI-to-text on Subject 1. Bold: best, underline: second best.
@@ 5 Conclusion
In this paper, we proposed "MindFormer", a powerful transformer architecture for semantically aligned multi-subject fMRI embedding for braining decoding. Thanks to the effective multi-subject fMRI signal embedding using the subject token and IP-Adapter, the model significantly outperformed the existing multi-subject brain decoding framework. Overall, MindFormer provides a new framework for understanding multi-subject brain decoding and common neural patterns. The model’s ability to leverage shared information across subjects while maintaining individual-specific accuracy marks a significant advancement in the field of brain decoding.
Limitation. Current implementation of MindFormer primarily focuses on visual stimuli. Extending this approach to decode more complex cognitive and sensory experiences will require substantial advancements in both model architecture and training methodologies. Another limitation is the computational complexity associated with training much more subjects. Although our approach reduces the parameter count compared to existing models, training these models with over about 10 subjects still require significant computational resources. Future work should aim to optimize the model further to make it more accessible and feasible for broader applications.
@@ A.1 Comparison of Model Parameter Number
Table 5: MindBridge and MindFormer have fewer parameters compared to other models. In particular, the single-MindFormer for one subject has the fewest number of parameters among them.
@@ A.2 Additional Results of multi MindFormer Model
Figure 9: Additional reconstructed image from human brain activity using MindFormer.
Figure 10: Additional reconstructed image from human brain activity using MindFormer.